OpenStudy (tabdo1226):

The Mendez family used 12 gallons of gasoline on a 300-mile trip. At that rate, how far would they go on 9 gallons of gasoline? mi

OpenStudy (tabdo1226):

@texaschic101

OpenStudy (tabdo1226):

i think it's 225 miles (mi)

OpenStudy (tabdo1226):

@Hotchellerae21

OpenStudy (texaschic101):

proportion time 12/300 = 9/x --- 12 gallons to 300 miles = 9 gallons to x miles cross multiply 12x = (300)(9) 12x = 2700 x = 2700/12 x = 225 so 9 gallons will get you 225 miles questions ?

OpenStudy (tabdo1226):

so im correct

OpenStudy (texaschic101):

yes...you are :) nice work...you deserve a medal

OpenStudy (tabdo1226):

you too and thanks

OpenStudy (texaschic101):

YW :)